40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 806     Entry time: Wed Aug 6 22:19:07 2008
Author: Yoichi 
Type: Update 
Category: SUS 
Subject: BS alignment 
Koji, Yoichi

We realized that we did not pay attention to the BS alignment while working on the alignment of the ITMX today. Because we were injecting the ALM laser (absolute length measurement laser) from the AS port, the ITMX alignment depends on the BS alignment.
The BS optical lever was not centered and the sum was about 2000cnt, which is low compared, for example, to the SRM oplev.
So we were not sure if the BS was in a good alignment or not.
So we decided to move the BS to center the QPD.
In doing so, we also moved the ITMX so that we do not lose the ALM laser beam coming back to the AS port.
When the BS oplev was centered, the sum of the QPD was still about 2000. So it was not far off centered.
After the tweaking, we were able to see some interference between the light reflected by the ITMY and ITMX at the AS port (actually this is the bright port for the ALM laser). By tweaking the ITMY, we were able to see Michelson fringes at the AS port.
If we believe the ALM laser alignment is still good after the vent, the ITMX, ITMY, BS and SRM should be now in a good alignment condition.
The OSEM values for the ITMX, BS, SRM seem to be ok (0.9+/-0.2). The ITMY LL is a bit low (~ 0.45).
ELOG V3.1.3-